OpenStudy (anonymous):

Totally having a brain fart here... can you walk me through solving the indefinite integral: sin(2x)+2

OpenStudy (zarkon):

\[\int\sin(ax)dx=\frac{-\cos(ax)}{a}+c\]

OpenStudy (anonymous):

Got it! Thanks

OpenStudy (zarkon):

good :)